Problem 1: Find sin(x) and cos(x) in the right triangle shown below. h 8 6
Answer #1

h=√(8^2+6^2)=√(64+36)=10

Therefore

Sin (x)=8/10 =4/5

And cos (x)=6/10=3/5

.